How much heat is needed to raise the temperature of 101 g of water from 23.0 °C to 92.0 °C?
Answer #1

This problem needs the concept of Calorimetry. I provide a detailed solution as follows, Hope it will help you.

7 Mass of water = 101 g Specific heat of water = leal/g.k Change of temperature = (92-23°c 69°c = (69+273) K. 342K so Heat Re